# Game Theory in Cosmology

### v1

@@ 8 Summary
We present a novel, potential game-theory-inspired statistical framework for cosmology, called Cosmological Teleodynamics. Our approach applies a maximum-caliber principle to cosmic histories, introducing a teleodynamic bias functional, $\Phi$, that encodes structural and evolutionary memory, yielding an additional drift force term $\alpha\nabla\Phi$ in the Boltzmann and Euler equations. In this framework, the dark sector, cosmic acceleration, large-scale structure formation, and the resolution of observational tensions emerge naturally from nonlocal memory effects and persistent self-organization in the gravitating Universe, rather than from new fundamental particles or fields.
@@
The bias functional decomposes into $\Phi=\bar{\Phi}+\varphi$, with $\bar{\Phi}$ generating an effective homogeneous energy density $\rho_{\text{TD}}$ that naturally drives late-time acceleration, thus explaining dark energy. The inhomogeneous component $\varphi$ modifies the Poisson equation through an emergent clustering density $\rho_{\text{TD}}(k,a)$, which reproduces dark matter-like behavior, including rotation curves, lensing, cluster dynamics, and scale-dependent growth. Different response kernels $K(k,a)$ allow teleodynamic “dark matter” to mimic CDM, WDM, SIDM, etc., as well as generate anisotropic, environment-dependent halo signatures aligned with the cosmic web. This framework can help alleviate current observational tensions, particularly the $H_{0}$ and $S_{8}$ discrepancies.
@@
It also predicts novel phenomena that cannot arise from conventional particle dark matter or scalar-field dark energy, including anisotropic velocity fields and environment-dependent halo signatures. Teleodynamic entropy emerges from microstate counting at the Hubble horizon, revealing a fundamentally nonequilibrium origin of cosmic entropy without invoking black hole thermodynamics. The current expansion epoch of the universe can be interpreted as it evolving towards an arbitrage equilibrium, in which it tries to maximize the global effective utility of accessible microstates in every region, thereby giving rise to the Law of Universal Arbitrage Equilibrium. The teleodynamic setup of cosmology can also support different fates for the universe, as different forms of the bias functional can lead to distinct long-term futures.
Cosmological Teleodynamics thus provides a unified, emergent, and observationally testable alternative to the standard dark sector paradigm. Rather than invoking new particles, it roots cosmological phenomena in the intrinsic statistical and systemic structure of cosmic memory, revealing that the Universe itself operates as a giant potential game.

### v2

@@ 10 Summary
We present a novel, potential game-theory-inspired statistical framework for cosmology, called Cosmic Teleodynamics. Our key insight is that, just as the Big Bang and Inflation (BBI) left their signature in the Cosmic Microwave Background (CMB), they also left similar indelible uneven structural features on spacetime, which continue to affect subsequent cosmic dynamics and evolution. We model this cosmic memory as a teleodynamic bias functional, $\Phi$, that encodes structural and evolutionary memory using the maximum-caliber principle, yielding an additional drift force term $\alpha\nabla\Phi$ in the Boltzmann and Euler equations. In this framework, the dark sector, cosmic acceleration, large-scale structure formation, and the resolution of observational tensions emerge naturally from nonlocal memory effects and persistent self-organization in the gravitating universe, rather than from new fundamental particles or fields.
@@
The bias functional decomposes into $\Phi=\bar{\Phi}+\varphi$, with $\bar{\Phi}$ generating an effective homogeneous energy density $\rho_{\text{TD}}$ that naturally drives late-time acceleration, thus explaining dark energy. The inhomogeneous component $\varphi$ modifies the Poisson equation through an emergent clustering density $\rho_{\text{TD}}(k,a)$, which reproduces dark matter-like behavior, including rotation curves, lensing, cluster dynamics, and scale-dependent growth. Different response kernels $K(k,a)$ allow teleodynamic "dark matter" to mimic CDM, WDM, SIDM, etc., as well as generate anisotropic, environment-dependent halo signatures aligned with the cosmic web. This framework can help alleviate current observational tensions, particularly the $H_{0}$ and $S_{8}$ discrepancies.
@@
It also predicts novel phenomena that cannot arise from conventional particle dark matter or scalar-field dark energy, including anisotropic velocity fields and environment-dependent halo signatures. Teleodynamic entropy emerges from microstate counting at the Hubble horizon, revealing a fundamentally nonequilibrium origin of cosmic entropy without invoking black hole thermodynamics. The current expansion epoch of the universe can be interpreted as it evolving towards an arbitrage equilibrium, in which it tries to maximize the global effective utility of accessible microstates in every region, thereby giving rise to the Law of Universal Arbitrage Equilibrium. We also showed that our theory addresses the coincidence problem in cosmology naturally through its entropy formalism for the cosmological horizon. The teleodynamic setup of cosmology can also support different fates for the universe, as different forms of the bias functional can lead to distinct long-term futures.
Since game theory is typically applied to living agents, such as in biology and economics, we wish to clarify that our work does not claim that the universe is "conscious" or that galaxies are agents in any intentional or cognitive sense. The use of game-theoretic language is purely mathematical and structural in the sense that potential games provide a mathematical framework for modeling systems that exhibit persistent, non-ergodic, bias-driven organization, regardless of whether the constituents possess goals or awareness. In our theory, galaxies belong to the third class of agents defined in statistical teleodynamics as intrinsically persistent systems whose behavior arises from physical constraints, such as long-range gravitational memory and accumulated structure, not from intention or cognition. The teleodynamic potential and effective “utility”, as we described in terms of microstates, etc., are simply mathematical tools to describe how such systems evolve toward attractor states when memory and nonlocal correlations are present. Thus, the framework treats galaxies the same way statistical mechanics treats molecules, but it is much more complete, combining statistical mechanics with the dynamics of goal-driven agents; no consciousness or agency is implied.
The use of game theoretic language in the present work is not intended to imply the existence of literal decision-making entities in the cosmological context. Rather, "agents" should be understood as coarse-grained elements of the gravitationally evolving system, such as halos, patches of the cosmic web, or effective phase-space regions, whose dynamics are governed by long-lived interactions and collective constraints. In this interpretation, the role of an agent is played by a macroscopic degree of freedom whose evolution is influenced not only by local conditions but also by the surrounding environment and this viewpoint is closely aligned with standard practices in statistical mechanics and dynamical systems, where collective behavior emerges from many-body interactions and is often captured through effective potentials or Lyapunov like functions rather than microscopic trajectories.
Within this coarse-grained perspective, the relevance of potential-game structures becomes clearer as potential games are characterized by the existence of a global scalar functional whose extremization governs the collective evolution of the system. In gravitational structure formation, similar organizing principles already appear implicitly, for example, in the tendency toward virialized states and attractor solutions. The teleodynamic bias functional plays an analogous role by encoding persistence and environmental feedback at the macroscopic level. Importantly, this does not introduce new forces or modify gravity at a fundamental level; instead, it provides a compact way to encode the statistical tendency of the system to evolve toward long-lived, dynamically stable configurations. The game theoretic analogy is therefore best viewed as a conceptual and mathematical organizing principle rather than a literal microphysical model. It motivates the introduction of a bias functional that captures collective persistence and environmental dependence, and it provides intuition for why such a functional can act as an effective potential governing large-scale evolution. The present work does not claim to derive cosmological structure formation from game theory but rather to show that a teleodynamic effective sector inspired by such ideas leads to a well-defined, covariant cosmological framework with novel and testable signatures. A more detailed exploration of the microscopic and mesoscopic origins of the bias functional, including its connection to statistical decision theory and non-equilibrium dynamics, will be considered in future work.
Cosmic teleodynamics thus offers a unified, emergent, and observationally testable alternative to the standard dark sector paradigm. Instead of proposing new particles, it connects cosmological phenomena to the intrinsic statistical and systemic structure of cosmic memory. Although these results are encouraging, much more work remains as we continue to explore. Our framework is just the beginning of a new direction in theoretical cosmology.